Find arc length of curve (sketch also) x=y^4 /8 + 1/4 y^2 , y=1 to y = 4


1
Expert's answer
2021-02-23T09:18:06-0500

Find arc length of curve (sketch also) "x=\frac{y^4}{8} + \frac{1}{4y^2}" , "y=1" to "y = 4" .

Arc length of curve:

"L=\int_{1}^{4} \sqrt{1+(\frac{dx}{dy})^2}dy=\int_{1}^{4} \sqrt{1+(\frac{y^3}{2}-\frac{1}{2y^3})^2}dy=\int_{1}^{4} \sqrt{(\frac{y^3}{2}+\frac{1}{2y^3})^2}dy=\int_{1}^{4} (\frac{y^3}{2}+\frac{1}{2y^3})dy="

"(\frac{y^4}{8}-\frac{1}{4y^2})|_1^4=\frac{256}{8}-\frac{1}{4\cdot 16}-(\frac18-\frac14)=32\frac{7}{64}"



Answer: "L=32\frac{7}{64}."